# A linear Convergence result for the Jacobi-Proximal Alternating Direction Method of Multipliers

### v2

@@
Abstract.
In this paper, we analyze the convergence rate of the Jacobi-Proximal Alternating Direction Method of Multipliers (ADMM), a method initially introduced by Deng et al.[6] for the block-structured optimization problem with linear constraint. We establish the linear convergence of the algorithm when the cost functions are strongly convex and smooth.
Key words and phrases:
Keywords: Alternating Direction Method of Multipliers, Jacobi-Proximal ADMM, Linear Convergence, Convex Optimization
1. Introduction
@@
$\begin{split}\min_{x_{1},\cdots,x_{N}}&\sum_{i=1}^{N}\tilde{f}_{i}(x_{i})\\ \textrm{subject to}~{}&\sum_{i=1}^{N}A_{i}x_{i}=c,\\ &x_{i}\in\mathcal{X}_{i},\qquad i=1,\cdots,N.\end{split}$ (1.1)
where $N\geq 2$, $x_{i}\in\mathbb{R}^{n_{i}},~{}A_{i}\in\mathbb{R}^{m\times n_{i}},~{}c\in\mathbb{R}^{m}~{}\textrm{and~{}}\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ are closed convex sets and $\tilde{f}_{i}:\mathbb{R}^{n_{i}}\to\left(-\infty,+\infty\right]~{}(i=1,2,\cdots,N)$ are closed proper convex functions. The solution set of (1.1) is assumed to be nonempty. Note that the convex constraints $x_{i}\in\mathcal{X}_{i}$ can be incorporated into the objective $\tilde{f}_{i}$ using the indicator function:
$I_{\mathcal{X}_{i}}(x_{i})=\left\{\begin{array}[]{ll}0&\textrm{if }x_{i}\in\mathcal{X}_{i},\\ +\infty&\textrm{otherwise.}\end{array}\right.$ (1.2)
Then the problem (1.1) can be rewritten as
@@
where $f_{i}(x_{i})=\tilde{f}_{i}(x_{i})+I_{\mathcal{X}_{i}}(x_{i})$. Thus, we consider the equivalent reformulation (1.1) throughout this paper for ease of presentation.
The problem (1.1) appears in various applications that contain the image alignment problem in [33], the robust principal component analysis model with noisy and incomplete data in [36], the latent variable Gaussian graphical model selection in [3, 31] and the quadratic discriminant analysis model in [30].
To solve the problem (1.1), various algorithms have been proposed. A simple distributed algorithm for solving (1.1) is dual decomposition [11], which is essentially a dual ascent method or dual subgradient method [35] as follows. Consider the Lagrangian for problem (1.1):
@@
where $\lambda\in\mathbb{R}^{m}$ is the Lagrangian multiplier or the dual variable. The dual decomposition updates as follows: for $k\geq 1$,
@@
and thus they can be carried out in parallel. With a suitable choice of $\alpha_{k}$ and certain assumptions, dual decomposition is guaranteed to converge to an optimal solution [35]. However, dual decomposition exhibits slow convergence in practice, with a convergence rate of $\mathcal{O}(1/\sqrt{k})$ for general convex problems.
An effective distributed algorithm for (1.1) can be designed based on the alternating direction method of multipliers (ADMM) [13, 15] using the following augmented Lagrangian:
@@
where $\rho>0$ is a penalty parameter. In the context of ADMM using the augmented Lagrangian, the primal variable $x_{i}$ can be updated using different approaches. One common method is the Gauss-Seidel update, in which the primal variable $x_{i}$ is updated sequentially as follows:
@@
Alternatively, the Jacobi-type update provides a parallel update scheme for the primal variable $x_{i}$ as follows:
@@
Since Gauss-Seidel ADMM updates the primal variables $x_{i}$ using the most recent information, the performance of Jacobi ADMM is typically expected to be worse than that of Gauss-Seidel ADMM. For instance, the work [19] found an example for which Jacobi ADMM diverges even when $N=2$. This issue was addressed by Deng et al.[6] that introduced the Jacobi-Proximal ADMM involving a proximal term $\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}$ for each $x_{i}$-subproblem with positive semi-definite matrix $P_{i}$ and a damping parameter $\gamma>0$ for the update of $\lambda$ to guarantee its convergence. The Jacobi-Proximal ADMM is formulated as in Algorithm 1.
@@
  Initialize: $x_{i}^{0}~{}(i=1,2,\cdots,N)$ and $\lambda^{0}$;
  for $k=0,1,\cdots$ do
     Update $x_{i}$ for $i=1,\cdots,N$ in parallel by:
@@
Update $\textrm{Update }\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg{(}\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg{)}.$ (1.6)
  end for
Here we denote $\|x_{i}\|^{2}_{P_{i}}:=x_{i}^{T}P_{i}x_{i}$. The Jacobi-Proximal ADMM has the advantage of being highly parallelizable, as the updates for $x_{i}$ for each $i=1,2,\ldots,N$ can be computed simultaneously for each $i$, offering potential computational efficiency gains. We note that the original $x_{i}$-subproblem (1.5) contains a quadratic term of the form $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$. If $A_{i}^{T}A_{i}$ is ill-conditioned or computationally expensive to invert, one approach is to modify $P_{i}$ as $P_{i}=D_{i}-\rho A_{i}^{T}A_{i}$, where $D_{i}$ is a well-conditioned and simple matrix(e.g., a diagonal matrix). This modification eliminates the quadratic term $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$, making the $x_{i}$-subproblem easier to solve. There are various choices for $P_{i}$ that can simplify the computation of the $x_{i}$-subproblem. Here are two widely used formulations for $P_{i}$ in [7], corresponding to the following methods:
  1. (1)
@@
The work of Deng et al.[6] established a sublinear convergence result for the convergence of Jacobi-Proximal ADMM when the proximal matrices $P_{i}$ satisfy $P_{i}\succ(\frac{N}{2-\gamma}-1)A_{i}^{T}A_{i}$ for $i=1,2,\ldots,N$ with $\gamma<2$. In particular, under specific choices of $P_{i}$, the convergence conditions reduce to $\tau_{i}>\rho(\frac{N}{2-\gamma}-1)\|A_{i}\|^{2}$ for the standard proximal method and $\tau_{i}>\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}$ for the prox-linear method.
We then extend the analysis to establish the linear convergence of Jacobi-Proximal ADMM. It is well-known that ADMM converges linearly when the cost functions exhibit additional structure, such as strong convexity. We prove that the Jacobi-Proximal ADMM converges linearly under the assumption that each cost function $f_{j}$ is strongly convex. We derive a condition on the parameters $\rho>0$ and $\gamma>0$ for linear convergence, and we establish a linear convergence rate dependent on these parameters. We refer to Section 2 for the details and we provide a comparison of our convergence result with previous work on the problem (1.1) in Table 1.
Ref. Convexity Regularity Rank Conditions Rate
Gauss-Seidel ADMM for $N=2$ [7] $f_{1}$: SC $f_{1}$: $L$-smooth $A_{1}$: full column rank $A_{2}$: full row rank linear
$f_{1},f_{2}$: SC $f_{1}$: $L$-smooth $A_{2}$: full row rank
$f_{1}$: SC $f_{1},f_{2}$: $L$-smooth $A_{1}$: full column rank
@@
Gauss-Seidel ADMM for $N\geq 3$ [26] $f_{2},\ldots,f_{N}$: SC $f_{N}$: $L$-smooth $A_{N}$: full row rank linear
$f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth
$f_{2},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $A_{1}$: full column rank
Gauss-Sediel ADMM for $N\geq 3$ [22] $f_{i}(x_{i})=g_{i}(E_{i}x_{i})+h_{i}(x_{i})$ $g_{i}$: strictly C $h_{i}$: C $g_{i}$: $L$-smooth $E_{i}$: full column rank linear
Jacobi-Proximal ADMM for $N\geq 3$ [6] $f_{1},\ldots,f_{N}$: C $A_{i}$: full column rank sub-linear
Jacobi-Proximal ADMM for $N\geq 3$ This work $f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth $\big{[}A_{1}^{T};~{}A_{2}^{T};~{}\cdots;~{}A_{N}^{T}\big{]}$: full row rank linear
Table 1. This table summarizes the convergence results for ADMM. In the ’Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex). In the ’Rate’ column, ’linear’ refers to the cases with linear convergence, while ’sublinear’ indicates the case with sublinear convergence. $N$ denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
1.1. Related works
The Alternating Direction Method of Multipliers (ADMM) [13, 15] is an effective approach that simplifies the original problem by decomposing it into multiple easier-to-solve minimization subproblems. Building on the classical Douglas-Rachford operator splitting method [8], Glowinski and Marrocco [16], along with Gabay and Mercier [14], introduced the well-known alternating direction method of multipliers (ADMM) with Gauss-Seidel decomposition for solving (1.1) when $N=2$. It utilizes the augmented Lagrangian for (1.1):
@@ 1.1. Related works
which incorporates a quadratic penalty of the constraints (with a parameter $\rho>0$) into the Lagrangian. At each iteration, the augmented Lagrangian is sequentially minimized with respect to $x_{1}$ and $x_{2}$, followed by an update of the dual variable $\lambda$. The ADMM has been extensively investigated for solving two-block convex minimization problems (i.e., $N=2$). Its global convergence in the case of $N=2$ has been established in [9, 12]. In a more recent work by Deng and Yin [7], a generalized version of ADMM was introduced, where proximal terms were added to the two subproblems in (2.2), demonstrating linear convergence under certain assumptions on the strong convexity of functions $f_{1}$ and $f_{2}$, as well as the rank conditions on $A_{1}$ and $A_{2}$. Additionally, various studies have explored the convergence rate properties of ADMM when $N=2$ (see, e.g., [28, 10, 18, 1, 29, 20, 32, 21]). Although the convergence properties of ADMM for $N=2$ are well understood, its behavior for $N\geq 3$ has remained uncertain for a considerable period. In the work by Chen et al. [4], a counterexample was provided, demonstrating that without additional conditions, the ADMM for $N\geq 3$ may fail to converge. Building on this, we first discuss studies that establishes the convergence conditions for ADMM using the Gauss-Seidel scheme when $N\geq 3$, where the Gauss-Seidel method updates variables sequentially. We refer to the works [2], [23], [5], and [27] that achieved global and sublinear convergence results of the 3-block ADMM. Han and Yuan [17] established the global convergence of ADMM (2.2) under the assumption that $f_{1},\cdots,f_{N}$ are all strongly convex and $\gamma$ is restricted to a specific range. Both Chen, Shen, and You [5] and Lin, Ma, and Zhang [25] relax the conditions compared to [17] by allowing only $N-1$ functions to be strongly convex, with with Chen et al. [5] considering the case $N=3$ and Lin et al. [25] considering $N\geq 3$. Furthermore, Lin, Ma, and Zhang [25] demonstrated that under these conditions, ADMM (2.2) exhibits a sublinear convergence rate in both the ergodic and non-ergodic senses. Lin, Ma, and Zhang [24] further presented sufficient conditions to guarantee a sublinear convergence rate for ADMM when $N\geq 3$. Lin, Ma, and Zhang [26] proved the global linear convergence of ADMM under the three scenarios. The global R-linear convergence of ADMM has been established by Hong and Luo [22]. Alternatively, Jacobi updates in ADMM have been extensively studied, offering parallelization advantages for multi-block optimization problems. The linear convergence of Gauss-Seidel ADMM, Proximal ADMM, and modified Jacobi ADMM, as shown in [22], is established under several assumptions. In [6], it was established that the multi-block Jacobi ADMM converges globally at a rate of $o(1/k)$ in two cases: (i) when all coefficient matrices are nearly mutually orthogonal and possess full column rank, or (ii) when proximal terms are introduced into the subproblems. In [19], the multi-block Jacobi ADMM was combined with a relaxation step, achieving a worst-case convergence rate of $O(1/k)$ in both ergodic and non-ergodic senses. In a recent work by Xiao, Deng, and Xu [38], the authors introduce the Distributed Jacobi-Proximal ADMM algorithm for solving consensus optimization problems.
1.2. Contribution

### v4

@@
Abstract
This paper is concerned with the Jacobi-Proximal Alternating Direction Method of Multipliers (Jacobi-Proximal ADMM) proposed in the works He, Xu, and Yuan (J. Sci. Comput., 2015) [On_the_proximal] and Deng et al. (J. Sci. Comput., 2017) [Parallel_multi-block], a parallel implementation of the ADMM for large-scale multi-block optimization problems with a linear constraint. While the $\mathit{o}(1/k)$ convergence of the Jacobi-Proximal ADMM for the case $N\geq 3$ has been well established in previous work, to the best of our knowledge, its linear convergence for $N\geq 3$ remains unproven. We establish the linear convergence of the algorithm when the cost functions are strongly convex and smooth. Numerical experiments are presented supporting the convergence result.
keywords:
Alternating Direction Method of Multipliers, Jacobi-Proximal ADMM, Linear Convergence, Convex Optimization
pacs:
MSC Classification]90C25
1 Introduction
@@
where $N\geq 2$, $x_{i}\in\mathbb{R}^{n_{i}},~A_{i}\in\mathbb{R}^{m\times n_{i}},~c\in\mathbb{R}^{m}$ and $f_{i}:\mathbb{R}^{n_{i}}\to\left(-\infty,+\infty\right]~(i=1,2,\cdots,N)$.
The problem (1.1) appears in various applications such as the image alignment problem in [robust_alignment], the robust principal component analysis model with noisy and incomplete data in [recovering_lowrank], the latent variable Gaussian graphical model selection in [latent_variable, node-based_learning] and the quadratic discriminant analysis model in [discriminant_analysis].
To solve the problem (1.1), various algorithms have been proposed. A simple distributed algorithm for solving (1.1) is dual decomposition [Generalized_Lagrange], which is essentially a dual ascent method or dual subgradient method [Minimization_methods]. To describe it, we consider the Lagrangian for problem (1.1):
@@
where $\lambda\in\mathbb{R}^{m}$ is the dual variable. Then, the dual decomposition updates as follows: for $k\geq 1$,
@@
and thus they can be carried out in parallel. With a suitable choice of $\alpha_{k}$ and certain assumptions, dual decomposition is guaranteed to converge to an optimal solution [Minimization_methods]. However, dual decomposition exhibits slow convergence in practice, with a convergence rate of $\mathit{O}(1/\sqrt{k})$ for general convex problems.
An effective distributed algorithm for (1.1) can be designed based on the alternating direction method of multipliers (ADMM) [A_dual, Sur_l] using the following augmented Lagrangian:
@@
where $\rho>0$ is a penalty parameter. In the context of ADMM using the augmented Lagrangian, the primal variable $x_{i}$ can be updated using different approaches. One common method is the Gauss-Seidel update, in which the primal variable $x_{i}$ is updated sequentially using the most recent values. This scheme is commonly referred to as the Gauss-Seidel ADMM, and is given by:
@@
The Gauss-Seidel ADMM was proposed by Wei and Ozdaglar [Distributed_alternating], where the authors established a convergence rate of $\mathit{O}(1/k)$ under the assumption that each objective function $f_{i}$ is convex. Alternatively, the Jacobi-type update, commonly referred to as the Jacobi ADMM, provides a parallel update scheme for the primal variable $x_{i}$ as follows:
@@
This scheme is suitable for parallel computation, and thus particularly effective for large-scale or big-data scenarios for the model (1.1). Deng et al. [Parallel_multi-block] established an $\mathit{o}(1/k)$ convergence rate for the Jacobi ADMM, assuming that the objective functions $f_{i}$ are convex and that the matrices $A_{i}$ are full column-rank and mutually near-orthogonal. However, with the same penalty parameter $\rho$, the Jacobi ADMM, which updates $x_{i}$ in parallel using previous-iterate information, typically exhibits inferior performance compared to the Gauss-Seidel ADMM, which updates blocks sequentially with the latest iterates. In particular, the work [On_full] found an example for which Jacobi ADMM diverges even when $N=2$. This issue was addressed by He, Xu, and Yuan [On_the_proximal], who proposed the Jacobi-Proximal ADMM. To ensure convergence, they introduced one key modification to the Jacobi ADMM:
  • Proximal term $\frac{s\rho}{2}\|A_{i}(x_{i}-x_{i}^{k})\|^{2}$: A proximal term $\frac{s\rho}{2}\|A_{i}(x_{i}-x_{i}^{k})\|^{2}$ was added to each $x_{i}$-subproblem, where $\rho>0$ is the penalty parameter and $s>0$ is the proximal coefficient that enforces the new iterate to remain sufficiently close to the previous one.
Deng et al. [Parallel_multi-block] later studied the algorithm in a slightly different parameter setting, which is essentially the same as that of He, Xu, and Yuan [On_the_proximal]. This variant incorporates the following two modifications to the Jacobi ADMM:
  • Proximal term $\mathbf{\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}}$: A proximal term $\frac{1}{2}\|x_{i}-x_{i}^{k}\|_{P_{i}}^{2}$ was added to each $x_{i}$-subproblem, where $P_{i}$ is a positive semi-definite matrix. This promotes strong convexity of the $x_{i}$-subproblems, thereby guaranteeing the uniqueness of their solutions even when the original objective functions are not strictly convex.
  • Damping parameter $\mathbf{\gamma>0}$: The dual update
    $\lambda^{k+1}=\lambda^{k}-\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg)$ (1.3)
    was modified to
    $\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg),$ (1.4)
    where $\gamma>0$ is a damping parameter used to control the step size and improve the algorithm’s stability.
The Jacobi-Proximal ADMM is formulated as in Algorithm 1.
@@
1:Initialize: $x_{i}^{0}\ (i=1,2,\ldots,N)$ and $\lambda^{0}$;
2:for $k=0,1,\ldots$ do
3:  Update $x_{i}$ for $i=1,\ldots,N$ in parallel by:
@@
$\textrm{Update }\lambda^{k+1}=\lambda^{k}-\gamma\rho\bigg(\sum_{i=1}^{N}A_{i}x_{i}^{k+1}-c\bigg).$ (1.6)
4:end for
Here we denote $\|x_{i}\|^{2}_{P_{i}}:=x_{i}^{T}P_{i}x_{i}$. The Jacobi-Proximal ADMM has the advantage of being highly parallelizable, as the updates for $x_{i}$ for each $i=1,2,\ldots,N$ can be computed simultaneously for each $i$, offering potential computational efficiency gains. We note that the original $x_{i}$-subproblem (1.5) contains a quadratic term of the form $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$. If $A_{i}^{T}A_{i}$ is ill-conditioned or computationally expensive to invert, one approach is to modify $P_{i}$ as $P_{i}=D_{i}-\rho A_{i}^{T}A_{i}$, where $D_{i}$ is a well-conditioned and simple matrix(e.g., a diagonal matrix). This modification eliminates the quadratic term $\frac{\rho}{2}x_{i}^{T}A_{i}^{T}A_{i}x_{i}$, making the $x_{i}$-subproblem easier to solve. There are various choices for $P_{i}$ that can simplify the computation of the $x_{i}$-subproblem. Here are two widely used formulations for $P_{i}$ in [On_the_global], corresponding to the following methods:
  1. 1.
@@
The work of Deng et al.[Parallel_multi-block] established an $\mathit{o}(1/k)$ convergence result for the convergence of the Jacobi-Proximal ADMM when the proximal matrices $P_{i}$ satisfy $P_{i}\succ(\frac{N}{2-\gamma}-1)A_{i}^{T}A_{i}$ for $i=1,2,\ldots,N$ with $\gamma<2$. In particular, under specific choices of $P_{i}$, the convergence conditions reduce to $\tau_{i}>\rho(\frac{N}{2-\gamma}-1)\|A_{i}\|^{2}$ for the standard proximal method and $\tau_{i}>\frac{\rho N}{2-\gamma}\|A_{i}\|^{2}$ for the prox-linear method.
In this work, we prove that the Jacobi-Proximal ADMM converges linearly under the assumption that each cost function $f_{j}$ is strongly convex. We derive a condition on the parameters $\rho>0$ and $\gamma>0$ for linear convergence, and we establish a linear convergence rate dependent on these parameters. We refer to Section 2 for the details and we provide a comparison of our convergence result with previous work on the problem (1.1) in Table 1.
Ref.
Convexity
Regularity
Rank
Conditions
Rate
Original ADMM
for $N=2$
[Local_linear]
$f_{1}(x_{1})+f_{2}(x_{2}$)
$\frac{1}{2}x_{1}^{T}Qx_{1}+q^{T}x_{1}+\frac{1}{2}x_{2}^{T}Rx_{2}+r^{T}x_{2}$
Local error bound
$B$: full column rank
$\mathit{O}(\sigma^{k})$(local)
Original ADMM
for $N=2$
[Linear_convergence_of_the]
$f_{1},f_{2}$: C
Global error bound
$A,B$: full column rank
$\mathit{O}(\sigma^{k})$
Original ADMM
for $N=2$
[Discerning_the]
Structured polyhedricity assumption
(see [Discerning_the, Assump. 1.3])
$\mathit{O}(\sigma^{k})$(local)
Gauss-Seidel
ADMM
for $N\geq 3$
[On_the_global_linear]
$f_{2},\ldots,f_{N}$: SC
$f_{N}$: $L$-smooth
$A_{N}$: full row rank
$\mathit{O}(\sigma^{k})$
$f_{1},\ldots,f_{N}$: SC $f_{1},\ldots,f_{N}$: $L$-smooth
$f_{2},\ldots,f_{N}$: SC
$f_{1},\ldots,f_{N}$: $L$-smooth
$A_{1}$: full column rank
Gauss-Seidel
ADMM
for $N\geq 3$
[On_the_linear]
$f_{i}(x_{i})=g_{i}(E_{i}x_{i})+h_{i}(x_{i})$
$g_{i}$: strictly C
$h_{i}$: C
$g_{i}$: $L$-smooth
$\operatorname{epi}(h_{i})$: polyhedral
$X_{i}$: compact polyhedral sets
Primal/Dual error bound
$E_{i}$: full column rank
$\mathit{O}(\sigma^{k})$
Jacobi
ADMM
for $N\geq 3$
[Parallel_multi-block]
$f_{1},\ldots,f_{N}$: C
$A_{i}$: mutually near-orthogonal
$\mathit{o}(1/k)$
Jacobi-Proximal
ADMM
for $N\geq 3$
[Parallel_multi-block]
$f_{1},\ldots,f_{N}$: C
$\mathit{o}(1/k)$
Jacobi-Proximal
ADMM
for $N\geq 3$
This
work
$f_{1},\ldots,f_{N}$: SC
$f_{1},\ldots,f_{N}$: $L$-smooth
$\big[A_{1}^{T};~A_{2}^{T};~\cdots;~A_{N}^{T}\big]$: full column rank
$\mathit{O}(\sigma^{k})$
Table 1: This table summarizes the convergence results for ADMM. In the ‘Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex). In the ‘Rate’ column, $\mathit{O}(\sigma^{k})$ denotes linear convergence with $0<\sigma<1$ ($\mathit{O}(\sigma^{k})$(local) indicates local linear convergence, and $\mathit{o}(1/k)$ indicates convergence strictly faster than sublinear rate. NN denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
Ref.
Convexity
Regularity
Rank
Conditions
Rate
Proximal
ADMM
[Partial_error]
$f_{1},f_{2}$:C
Strong Conical Hull
Intersection Property
If $P_{1}\succeq 0$: $A_{1},A_{2}$: full column rank
If $P_{1}\succ 0$: $A_{2}$: full column rank
$\mathit{O}(\sigma^{k})$
PADMM-FG
[On_the_global]
$f_{1}$: SC
$f_{1}$: $L$-smooth
$A_{1}$: full column rank
$A_{2}$: full row rank
$\mathit{O}(\sigma^{k})$
$f_{1},f_{2}$: SC
$f_{1}$: $L$-smooth
$A_{2}$: full row rank
$f_{1}$: SC $f_{1},f_{2}$: $L$-smooth
$A_{1}$: full column rank
@@
PADMM-FG
[Discerning_the]
Structured polyhedricity assumption
(see [Discerning_the, Assump. 1.3])
$\mathit{O}(\sigma^{k})$(local)
Structured subregularity assumption (see [Discerning_the, Assump. 1.4])
$A_{1}$: full column rank
$A_{2}$: full row rank
Relaxed
ADMM
[metric_selection]
$f_{1}$:SC
$f_{1}$:$L$-smooth
AA: full row rank
$\mathit{O}(\sigma^{k})$
Relaxed
ADMM
[Faster_convergence]
$f_{1}$: SC
$f_{1}$: $L$-smooth
$A_{1}$: full row rank
$\mathit{O}(e^{-\alpha k})$
$f_{2}$: SC
$f_{2}$: $L$-smooth
$A_{2}$: full row rank
$f_{2}$: SC $f_{1}$: $L$-smooth
$A_{1}$: full row rank
$f_{1}$: SC $f_{2}$: $L$-smooth
$A_{2}$: full row rank
Over-Relaxed
ADMM
[A_general]
$f_{1}$:SC, $f_{2}$:C
$f_{1}$: $L$-smooth
$A_{2}$: full column rank
$\mathit{O}(\sigma^{k})$
Linearized
ADMM
[Discerning_the]
Structured polyhedricity assumption
(see [Discerning_the, Assump. 1.3])
$\mathit{O}(\sigma^{k})$(local)
Structured subregularity assumption (see [Discerning_the, Assump. 1.4])
$A_{2}$: full row rank
Table 2: This table summarizes the convergence results for several variants of ADMM in the case $N=2$. Here, PADMM-FG refers to the proximal ADMM with the Fortin–Glowinski relaxation step. In the ‘Convexity’ column, C (resp., SC) means that the total cost function is assumed to be convex (resp., strongly convex). In the ‘Rate’ column, $\mathit{O}(\sigma^{k})$ with $0<\sigma<1$ and $\mathit{O}(e^{-\alpha k})$ with $\alpha>0$ both denote linear convergence, while $\mathit{O}(\sigma^{k})$(local) indicates local linear convergence. NN denotes the number of blocks of objective functions and $k$ denotes the number of iterations.
1.1 Related works
The Alternating Direction Method of Multipliers (ADMM) [A_dual, Sur_l] is an efficient approach that facilitates the solution of structured optimization problems by decomposing them into a sequence of simpler subproblems. Building on the classical Douglas–Rachford operator splitting method [On_the], Glowinski and Marrocco [Sur_l], along with Gabay and Mercier [A_dual], introduced the well-known Alternating Direction Method of Multipliers (ADMM) with a Gauss–Seidel-type decomposition for solving (1.1) in the case of $N=2$. It employs the augmented Lagrangian for (1.1):
@@ 1.1 Related works
which incorporates a quadratic penalty term of the constraints (with a penalty parameter $\rho>0$) into the Lagrangian. At each iteration, the augmented Lagrangian is sequentially minimized with respect to the primal variables $x_{1}$ and $x_{2}$, followed by an update of the dual variable $\lambda$. The ADMM has been extensively investigated for solving two-block convex minimization problems (i.e., $N=2$). Its global convergence in the case of $N=2$ has been established in [On_the_Douglas, Chapter_ix]. Han and Yuan [Local_linear] showed that original ADMM achieves a local linear convergence rate for quadratic programs by using an error bound established in [Error_bound]. Based on this error bound framework, Yang and Han [Linear_convergence_of_the] proved that ADMM attains global linear convergence for convex problems with polyhedral constraints by leveraging a global error bound. Yuan, Zeng, and Zhang [Discerning_the] showed that original ADMM also enjoys global linear convergence under the structured polyhedrality assumption by employing variational analysis techniques. These scenarios are summarized in Table 1. Additionally, various studies have explored the convergence rate properties of original ADMM when $N=2$ (see, e.g., [Splitting_algorithms, Local_linear_linear, On_the_O(1/n), Iteration-complexity_of, On_non-ergodic]). In addition to these results on the original ADMM, convergence of several variants for the case of $N=2$ has also been studied, such as the Proximal ADMM [Partial_error], Proximal ADMM with the Fortin–Glowinski relaxation step (PADMM-FG) [On_the_global, Discerning_the], Relaxed ADMM [metric_selection, Faster_convergence], Over-Relaxed ADMM [A_general], and Linearized ADMM [Discerning_the]. The convergence results of these variants are summarized in Table 2.
While the convergence results of ADMM is well established for the case $N=2$, its behavior for $N\geq 3$ has long remained unresolved; particularly, Chen et al. [The_direct] provided a concrete example in which Gauss-Seidel ADMM diverges for $N=3$ for any penalty parameter $\rho$. Building on this, we first discuss studies that establishes the convergence conditions for ADMM using the Gauss-Seidel scheme when $N\geq 3$, where the Gauss-Seidel method updates variables sequentially. We refer to the works [A_convergent], [convergence_analysis], and [Global_convergence] that achieved global and sublinear convergence results of the 3-block ADMM. Han and Yuan [A_note] established the global convergence of ADMM under the assumption that $f_{1},\cdots,f_{N}$ are all strongly convex. Building on this, both Chen, Shen, and You [convergence_analysis] and Lin, Ma, and Zhang [On_the_sublinear] relax the assumptions in [A_note]. In particular, [convergence_analysis], for $N=3$, allows one objective function to be not necessarily strongly convex and also relax the range of the penalty parameter, and propose a relaxed ADMM that incorporates an additional computation of optimal step size and prove its global convergence. Lin, Ma, and Zhang [On_the_sublinear] demonstrate sublinear convergence with rate $\textit{O}(1/k)$ in the ergodic sense and $\textit{o}(1/k)$ in the non-ergodic sense when one objective function is convex and the remaining N1N-1 are strongly convex, provided that the penalty parameter lies within a certain range. Lin, Ma, and Zhang [On_the_global_linear] proved the global linear convergence of Gauss-Seidel ADMM under the three scenarios, which are detailed in Table 1. The global R-linear convergence of Gauss-Seidel ADMM has been established by Hong and Luo [On_the_linear]; the corresponding condition is summarized in Table 1.
Alternatively, Jacobi updates in ADMM have been extensively studied, offering parallelization advantages for multi-block optimization problems. The linear convergence of Proximal ADMM and modified Jacobi ADMM, as well as that of Gauss-Seidel ADMM, which was mentioned earlier, is established under several assumptions in [On_the_linear]. Deng et al. [Parallel_multi-block] showed that the Jacobi ADMM and Jacobi-Proximal ADMM achieve $\mathit{o}(1/k)$ convergence rate under the conditions summarized in Table 1. He, Hou, and Yuan [On_full] proposed a variant of multi-block Jacobi ADMM by incorporating an under-relaxation step into the full Jacobian decomposition framework, achieving a worst-case convergence rate of $O(1/k)$ in both ergodic and non-ergodic senses.
1.2 Contribution